OpenStudy (anonymous):

what is the function rule of(0,1), (1,3), (2,9), (3,27), (4,81)

OpenStudy (anonymous):

y=2x+1 y=mx+b b=y-intercept which is (0,y) m=slope slope formula is: y2-y1/x2-x1

OpenStudy (anonymous):

thank you but thats not what im looking for im looking for somthing like y=2x - 2

OpenStudy (anonymous):

(0,1) is the y-intercept, -2 is not your y-intercept based on the points you given.

OpenStudy (jdoe0001):

\(\large \begin{array}{lrlll} x&y\\ \hline\\ 0&1\implies 3^0\\ 1&3\implies 3^1\\ 2&9\implies 3^2\\ 3&27\implies 3^3\\ 4&81\implies 3^4 \end{array}\) notice anything?

OpenStudy (anonymous):

that helps but i need the answer

OpenStudy (anonymous):

Maybe something like\[f(x)=3^x\]That function is defined for all values of x; based on the way it was described, the domain might only be\[x \in \left\{ 0.1.2.3.4 \right\}\]